\begin{thebibliography}{10}
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\expandafter\ifx\csname href\endcsname\relax
  \def\href#1#2{#2} \def\path#1{#1}\fi

\bibitem{kulyabov:2017:sfm:instantaneous_hamiltonian}
D.~S. Kulyabov, A.~V. Korolkova, L.~A. Sevastianov, E.~G. Eferina, T.~R.
  Velieva, I.~S. Zaryadov, {Maxwell's Equations Instantaneous Hamiltonian}, in:
  V.~L. Derbov, D.~E. Postnov (Eds.), Proceedings of SPIE. Saratov Fall Meeting
  2016: Laser Physics and Photonics XVII and Computational Biophysics and
  Analysis of Biomedical Data III, Vol. 10337 of Proceedings of SPIE - The
  International Society for Optical Engineering, SPIE, 2017, pp. 103370L.1--9.
\newblock \href {https://doi.org/10.1117/12.2267938}
  {\path{doi:10.1117/12.2267938}}.

\bibitem{kulyabov:2017:sfm:geometric_lagrangian}
D.~S. Kulyabov, A.~V. Korolkova, L.~A. Sevastianov, E.~G. Eferina, T.~R.
  Velieva, {A Geometric Approach to the Lagrangian and Hamiltonian Formalism of
  Electrodynamics}, in: V.~L. Derbov, D.~E. Postnov (Eds.), Proceedings of
  SPIE. Saratov Fall Meeting 2016: Laser Physics and Photonics XVII and
  Computational Biophysics and Analysis of Biomedical Data III, Vol. 10337 of
  Proceedings of SPIE - The International Society for Optical Engineering,
  SPIE, 2017, pp. 103370M.1--6.
\newblock \href {https://doi.org/10.1117/12.2267944}
  {\path{doi:10.1117/12.2267944}}.

\bibitem{sivukhin:1979:ufn::en}
D.~V. Sivukhin, {The international system of physical units}, Soviet Physics
  Uspekhi 22~(10) (1979) 834--836.
\newblock \href {https://doi.org/10.1070/PU1979v022n10ABEH005711}
  {\path{doi:10.1070/PU1979v022n10ABEH005711}}.

\bibitem{vilasi:book:hamiltonian::en}
G.~Vilasi, {Hamiltonian Dynamics}, World Scientific Pub Co Inc, 2001.

\bibitem{dirac:lectures-quantum-mechanics::en}
P.~A.~M. Dirac, {Lectures on Quantum Mechanics}, Courier Corporation, 2001.

\bibitem{bergman:1955:lagrangians}
E.~Newman, P.~G. Bergmann, {Lagrangians Linear in the "Velocities"}, Physical
  Review 99~(2) (1955) 587--592.
\newblock \href {https://doi.org/10.1103/PhysRev.99.587}
  {\path{doi:10.1103/PhysRev.99.587}}.

\bibitem{sardanashvily:1995:hamiltonian_formalism}
G.~A. Sardanashvily, {Generalized Hamiltonian Formalism for Field Theory},
  World Scientific Publishing Company, Singapore, 1995.
\newblock \href {https://doi.org/10.1088/0264-9381/13/12/025}
  {\path{doi:10.1088/0264-9381/13/12/025}}.

\bibitem{giachetta:1999:covariant_hamilton}
G.~Giachetta, L.~Mangiarotti, G.~A. Sardanashvily, {Covariant Hamilton
  Equations for Field Theory}, J Phys A 32~(38) (1999) 6629--6642.
\newblock \href {https://doi.org/10.1088/0305-4470/32/38/302}
  {\path{doi:10.1088/0305-4470/32/38/302}}.

\bibitem{giachetta:2009:advanced_classical}
G.~Giachetta, L.~Mangiarotti, G.~A. Sardanashvily, {Advanced Classical Field
  Theory}, World Scientific Publishing Company, Singapore, 2009.
\newblock \href {https://doi.org/10.1142/7189} {\path{doi:10.1142/7189}}.

\bibitem{stratton:1948::en}
J.~A. Stratton, {Electromagnetic Theory}, MGH, 1941.

\end{thebibliography}


\begin{thebibliography}{10}
\def\selectlanguageifdefined#1{
\expandafter\ifx\csname date#1\endcsname\relax
\else\selectlanguage{#1}\fi}
\providecommand*{\href}[2]{{\small #2}}
\providecommand*{\url}[1]{{\small #1}}
\providecommand*{\BibUrl}[1]{\url{#1}}
\providecommand{\BibAnnote}[1]{}
\providecommand*{\BibEmph}[1]{#1}
\ProvideTextCommandDefault{\cyrdash}{\iflanguage{russian}{\hbox
  to.8em{--\hss--}}{\textemdash}}
\providecommand*{\BibDash}{\ifdim\lastskip>0pt\unskip\nobreak\hskip.2em plus
  0.1em\fi
\cyrdash\hskip.2em plus 0.1em\ignorespaces}
\renewcommand{\newblock}{\ignorespaces}

\bibitem{kulyabov:2017:sfm:instantaneous_hamiltonian}
\selectlanguageifdefined{english}
\BibEmph{Kulyabov~D.~S., Korolkova~A.~V., Sevastianov~L.~A., Eferina~E.~G.,
  Velieva~T.~R., Zaryadov~I.~S.}
  \href{http://dx.doi.org/10.1117/12.2267938}{{Maxwell's Equations
  Instantaneous Hamiltonian}}~// Proceedings of SPIE. Saratov Fall Meeting
  2016: Laser Physics and Photonics XVII and Computational Biophysics and
  Analysis of Biomedical Data III~/ Ed.\ by\ V.~L.~Derbov, D.~E.~Postnov.
  \BibDash
\newblock Vol.~10337 of \BibEmph{Proceedings of SPIE - The International
  Society for Optical Engineering}. \BibDash
\newblock SPIE, 2017. \BibDash
\newblock P.~103370L.1--9.

\bibitem{kulyabov:2017:sfm:geometric_lagrangian}
\selectlanguageifdefined{english}
\BibEmph{Kulyabov~D.~S., Korolkova~A.~V., Sevastianov~L.~A., Eferina~E.~G.,
  Velieva~T.~R.} \href{http://dx.doi.org/10.1117/12.2267944}{{A Geometric
  Approach to the Lagrangian and Hamiltonian Formalism of Electrodynamics}}~//
  Proceedings of SPIE. Saratov Fall Meeting 2016: Laser Physics and Photonics
  XVII and Computational Biophysics and Analysis of Biomedical Data III~/ Ed.\
  by\ V.~L.~Derbov, D.~E.~Postnov. \BibDash
\newblock Vol.~10337 of \BibEmph{Proceedings of SPIE - The International
  Society for Optical Engineering}. \BibDash
\newblock SPIE, 2017. \BibDash
\newblock P.~103370M.1--6.

\bibitem{sivukhin:1979:ufn::ru}
\selectlanguageifdefined{russian}
\BibEmph{Сивухин~Д.~В.} {О Международной системе
  физических величин}~//
  \href{http://dx.doi.org/10.3367/UFNr.0129.197910h.0335}{\BibEmph{Успехи
  физических наук}}. \BibDash
\newblock 1979. \BibDash
\newblock Т. 129, {№}~10. \BibDash
\newblock {С.}~335--338.

\bibitem{vilasi:book:hamiltonian::ru}
\selectlanguageifdefined{russian}
\BibEmph{Вилази~Г.} {Гамильтонова динамика}.
  \BibDash
\newblock М.~: РХД, 2006. \BibDash
\newblock 432~{с.}

\bibitem{dirac:lectures-quantum-mechanics::ru}
\selectlanguageifdefined{russian}
\BibEmph{Дирак~П. А.~М.} {Лекции по квантовой
  механике}. \BibDash
\newblock Ижевск~: РХД, 1998.

\bibitem{bergman:1955:lagrangians}
\selectlanguageifdefined{english}
\BibEmph{Newman~E., Bergmann~P.~G.} {Lagrangians Linear in the "Velocities"}~//
  \href{http://dx.doi.org/10.1103/PhysRev.99.587}{\BibEmph{Physical Review}}.
  \BibDash
\newblock 1955. \BibDash jul. \BibDash
\newblock Vol.~99, no.~2. \BibDash
\newblock P.~587--592.

\bibitem{sardanashvily:1995:hamiltonian_formalism}
\selectlanguageifdefined{english}
\BibEmph{Sardanashvily~G.~A.}
  \href{http://dx.doi.org/10.1088/0264-9381/13/12/025}{{Generalized Hamiltonian
  Formalism for Field Theory}}. \BibDash
\newblock Singapore~: World Scientific Publishing Company, 1995. \BibDash
\newblock viii+155~p.

\bibitem{giachetta:1999:covariant_hamilton}
\selectlanguageifdefined{english}
\BibEmph{Giachetta~G., Mangiarotti~L., Sardanashvily~G.~A.} {Covariant Hamilton
  Equations for Field Theory}~//
  \href{http://dx.doi.org/10.1088/0305-4470/32/38/302}{\BibEmph{J Phys A}}.
  \BibDash
\newblock 1999. \BibDash sep. \BibDash
\newblock Vol.~32, no.~38. \BibDash
\newblock P.~6629--6642.

\bibitem{giachetta:2009:advanced_classical}
\selectlanguageifdefined{english}
\BibEmph{Giachetta~G., Mangiarotti~L., Sardanashvily~G.~A.}
  \href{http://dx.doi.org/10.1142/7189}{{Advanced Classical Field Theory}}.
  \BibDash
\newblock Singapore~: World Scientific Publishing Company, 2009. \BibDash
\newblock 382~p.

\bibitem{stratton:1948::ru}
\selectlanguageifdefined{russian}
\BibEmph{Стрэттон~Д.~А.} {Теория
  электромагнетизма}. \BibDash
\newblock М.-Л.: ГИТТЛ, 1948.

\end{thebibliography}

\documentclass[%
floatfix,
showkeys,
nofootinbib, %
superscriptaddress, %
]{revtex4-1}

\usepackage{cmap}

\usepackage{ucs}
\usepackage[utf8x]{inputenc}
\usepackage[T1,T2A]{fontenc}
\usepackage[german,russian,english]{babel}

\usepackage[sort&compress]{natbib}
\usepackage{amsmath}
\usepackage{amssymb}

\usepackage{mathtools}
\mathtoolsset{
showonlyrefs,
mathic = true
}

\allowdisplaybreaks

\usepackage{hyperref}
\hypersetup{backref,
 colorlinks=false}
\hypersetup{pdfborder=0 0 0}

\usepackage{microtype}
\UseMicrotypeSet[protrusion]{alltext}

\usepackage{graphicx}

\usepackage[scanall]{psfrag}

\usepackage{listings}
\usepackage{listingsutf8}
\usepackage{fixltx2e}

\usepackage{nicefrac}

\makeatletter
\def\ps@pprintTitle{%
     \let\@oddhead\@empty
     \let\@evenhead\@empty
     \let\@oddfoot\@empty
     \let\@evenfoot\@oddfoot}
\makeatother

\usepackage{physics}

\begin{document}

\title{Constrained Hamiltonian approach to the Maxwell theory}

\author{Dmitry S. Kulyabov}
\email{kulyabov-ds@rudn.ru}
\affiliation{Department of Applied Probability and Informatics,\\
  Peoples' Friendship University of Russia (RUDN University),\\
  6 Miklukho-Maklaya St, Moscow, 117198, Russian Federation}
\affiliation{Laboratory of Information Technologies\\
  Joint Institute for Nuclear Research\\
  6 Joliot-Curie, Dubna, Moscow region, 141980, Russia}

\author{Anna V. Korolkova}
\email{korolkova-av@rudn.ru}
\affiliation{Department of Applied Probability and Informatics,\\
  Peoples' Friendship University of Russia (RUDN University),\\
  6 Miklukho-Maklaya St, Moscow, 117198, Russian Federation}

\author{Migran N. Gevorkyan}
\email{gevorkyan-mn@rudn.ru}
\affiliation{Department of Applied Probability and Informatics,\\
  Peoples' Friendship University of Russia (RUDN University),\\
  6 Miklukho-Maklaya St, Moscow, 117198, Russian Federation}

\author{Leonid A. Sevastianov}
\email{sevastianov-la@rudn.ru}
\affiliation{Department of Applied Probability and Informatics,\\
  Peoples' Friendship University of Russia (RUDN University),\\
  6 Miklukho-Maklaya St, Moscow, 117198, Russian Federation}
\affiliation{Bogoliubov Laboratory of Theoretical Physics\\
  Joint Institute for Nuclear Research\\
  6 Joliot-Curie, Dubna, Moscow region, 141980, Russia}

\begin{abstract}
The most common physical formalisms are the Lagrangian formalism and
the Hamiltonian formalism. From the superficial point of view, they are
one and the same, but rewritten in other terms. However, it seems that
the Hamiltonian formalism has a richer structure and is more
convenient for studying the electromagnetic field, especially in the
formalization of its geometrization. Unfortunately for field problems,
there is a whole set of Hamiltonian formalisms. The authors study the
applicability of different variants of the Hamiltonian formalism to
the problems of electrodynamics. In this paper we consider the
Hamiltonian formalism with constraints.
\end{abstract}

  \keywords{Lagrangian formalism, Hamiltonian formalism, Hamiltonian formalism with constraints, Maxwell equations}

\maketitle

\section{Introduction}
\label{sec:intro}

In the study of electromagnetic and optical phenomena Hamiltonian formalism is often used. The main drawback of Hamiltonian formalism
it seems to be poorly developed for field systems. For
descriptions of field systems, we propose several variants of the Hamiltonian
formalism. Previously, we considered the possibility of building
symplectic~\cite{kulyabov:2017:sfm:instantaneous_hamiltonian} and
multipulse~\cite{kulyabov:2017:sfm:geometric_lagrangian}
Hamiltonian formalism. In this paper we consider the construction of a Hamiltonian formalism with constraints.

\section{Notations and conventions}
\label{sec:notation}

  \begin{enumerate}

  \item We will adhere to the following agreements. Greek indices
    ($\alpha$, $\beta$) will refer to the four-dimensional space.
    Latin
    indices from the middle of the alphabet ($i$, $j$, $k$) will refer
    to the three-dimensional space.

  \item In the theoretical description, Latin indices will refer to
    the space of arbitrary dimension.

  \item The comma in the index denotes a partial derivative with respect to
    corresponding coordinate ($f_{, i} := \partial_{i}f$);  the semicolon
    denotes a covariant derivative ($f_{;i} := \nabla_{i} f$).

  \item The CGS symmetrical system~\cite{sivukhin:1979:ufn::en} is used for
    notation of the equations of electrodynamics.

  \end{enumerate}

\section{Hamiltonian formalism}
\label{sec:hamilton}

There are several variants of the Hamiltonian formalism.
\begin{itemize}
\item The symplectic Hamiltonian formalism~\cite{vilasi:book:hamiltonian::en,kulyabov:2017:sfm:instantaneous_hamiltonian}.
\item The Dirac–Bergman Hamiltonian formalism for systems with constraints~\cite{dirac:lectures-quantum-mechanics::en, bergman:1955:lagrangians}.
\item The Hamilton–De Donder Hamiltonian formalism~\cite{sardanashvily:1995:hamiltonian_formalism}.
\item The multimomentum Hamiltonian formalism~\cite{sardanashvily:1995:hamiltonian_formalism,
    giachetta:1999:covariant_hamilton, giachetta:2009:advanced_classical}.
\end{itemize}

Consider major points of the Hamiltonian formalism.

Let the system be described by some quantity called the action:
\begin{equation}
S[q^i] = \int \dd^4 \mathcal{L}(x^i, q^i, \Dot{q}^i).
\end{equation}
Lagrangian (Lagrangian density)
$(x^I, q^i, \Dot{q}^i)$ depends on the generalized coordinates $q$ and their
first derivatives of $\Dot{q}$.

In the transition to the Hamiltonian formalism
the system is described by generalized coordinates
$q^I$ and generalized momentum:
\begin{equation}
  \label{eq:hamilton:p-L}
p_i = \pdv{\mathcal{L}}{\Dot{q}^i}.
\end{equation}

One can construct the Legendre transform in velocities:
\begin{equation}
\mathcal{H}(q^i, p_i) := p_i \Dot{q}^i − \mathcal{L}(q^i, \Dot{q}^i).
\end{equation}
The function $\mathcal{H}$ is called a Hamiltonian (Hamiltonian
density). The Hamiltonian of the system depends only on
generalized coordinates and momenta.

Using the connection between Hamiltonian and Lagrangian, we consider the action:
\begin{equation}
S[q^i,p_i] = \int \dd^4 \qty(p_i\Dot{q}^i - \mathcal{H}(q^i, p_i)).
\end{equation}
The corresponding system of Euler--Lagrange equations has the form:
\begin{equation}
  \label{eq:hamilton:ham-eq-S}
  \begin{gathered}
    \fdv{S}{q^i} = - \Dot{p}_i - \fdv{\mathcal{H}}{q^i} = 0,
    \\
    \fdv{S}{p_i} = \Dot{q}^i - \fdv{\mathcal{H}}{p_i} = 0.
  \end{gathered}
\end{equation}

Let us define the Poisson brackets:
\begin{equation}
[f, g] := \pdv{f}{q^i} \pdv{g}{p_i} - \pdv{f}{p_i} \pdv{g}{q^i}.
\end{equation}

Then we can rewrite equation~\eqref{eq:hamilton:ham-eq-S} as:
\begin{equation}
  \begin{gathered}
    \Dot{q}^i = \qty[q^i, \mathcal{H}] = \fdv{\mathcal{H}}{p_i},
    \\
    \Dot{p}_i = \qty[p_i, \mathcal{H}] = - \fdv{\mathcal{H}}{q^i}.
  \end{gathered}
\end{equation}

\section{Hamiltonian dynamics with constraints}
\label{sec:constrain}

If the Lagrangian is singular by velocities:
\begin{equation}
\det \qty[\pdv{\mathcal{L}}{\Dot{q}^i}{\Dot{q}^j}] = 0,
\end{equation}
it is not possible to express all momentum by
formula~\eqref{eq:hamilton:p-L}. In this case
one gets only possible momentum and for the rest
the concept of constraints~\cite{dirac:lectures-quantum-mechanics::en} is used.

We will consider the system with Lagrangian
$\mathcal{L} (x^k,q^k,p_k)$, $I=\overline{1,n}$. Also consider
set of constraints:
\begin{equation}
  \varphi^a (x^k,q^k,\Dot{q}^k), \quad a=\overline{1,m},
  \quad m \leqslant n.
\end{equation}

Action minimum, in case the trajectory satisfy the equations of connection,
is interpret also on trajectories without constraints, but with Lagrangian with
constraints: 
\begin{equation}
  L(x^k,q^k,\Dot{q}^k) =
  \mathcal{L}(x^k,q^k,\Dot{q}^k) -
  \lambda_a (x^k,q^k,\Dot{q}^k) \varphi^a (x^k,q^k,\Dot{q}^k).
\end{equation}

Hamilton's equations take the following form:
\begin{equation}
  \begin{gathered}
    \Dot{q}^i = \fdv{\mathcal{H}}{p_i} + \lambda_a \pdv{\varphi^a}{p_i},
    \\
    \Dot{p}_i = - \fdv{\mathcal{H}}{q^i} - \lambda_a \pdv{\varphi^a}{q^i}.
  \end{gathered}
\end{equation}

Lagrange multipliers are found from the condition of preserving constraints:
\begin{equation}
\Dot{\varphi^a} = [\varphi^a, \mathcal{H}] + \lambda_a \varphi^a = 0.
\end{equation}

\section{The Hamiltonian of the electromagnetic field with constraints}
\label{sec:em-hamilton}

We consider the construction of a Hamiltonian formalism with constraints for the case
electromagnetic field.

Write the Lagrangian of the electromagnetic field~\cite{stratton:1948::en}:
\begin{equation}
\mathcal{L} \qty(x^{\alpha}, A_{\beta}, A_{\alpha, \beta}) 
= 
- \frac{1}{16\pi c} F_{\alpha\beta} F^{\alpha\beta} 
\sqrt{-g}
-
\frac{1}{c^2} A_\alpha j^\alpha 
\sqrt{-g}.
\end{equation}

Since $F_{00}=0$,
\begin{equation}
\pdv[2]{\mathcal{L}}{\Dot{A}_0}=0.
\end{equation}
That is, the Lagrangian is irregular. Therefore, we need to find
constraints for the case $p^0$. Because
\begin{equation}
  p^0 = \pdv{\mathcal{L}}{\Dot{A}_0} = 0,
\end{equation}
then enter the constraints
\begin{equation}
  \label{eq:em-hamilton:varphi}
  \varphi := p^0 \approx 0.
\end{equation}
The $\approx$ symbol indicates that the equality must be performed on
surfaces of all constraints.

Let's write down expressions for momentum:
\begin{equation}
  p_i = \pdv{\mathcal{L}}{\Dot{A}^i} =
  \frac{1}{c} \pdv{\mathcal{L}}{A^i_{,0}}
  =
  -
  \frac{\sqrt{-g}}{4\pi c^2}
  \qty[
  A_{i,0} - A_{0,i}
  ]
  .
\end{equation}

Let us express time derivatives of $A_i$ by momentum:
\begin{equation}
\Dot{A}_i = c A_{i,0} = - \frac{4\pi c^3}{\sqrt{-g}} p_i + c A_{0,i}.
\end{equation}

Let us construct the Hamiltonian
\begin{equation}
  H =
  p^i \Dot{A}_i - \mathcal{L}
  +
  \lambda p^0
  =
  p^{i}\qty(-\frac{4\pi c^3}{\sqrt{-g}} p_i + c A_{0,i})
  +
  \frac{\sqrt{-g}}{16\pi c} F_{\alpha\beta} F^{\alpha\beta} 
  +
  \frac{\sqrt{-g}}{c^2} A_\alpha j^\alpha
  +
  \lambda p^0
\end{equation}

Hamilton's equations take the following form:
\begin{gather}
  \Dot{A}_0 = \pdv{H}{p^0} = \lambda,
  \label{eq:em-hamilton:dotA0}
  \\
  \Dot{A}_{i} = \pdv{H}{p^{i}} = 
  -
  \frac{4\pi c^3}{\sqrt{-g}} p_i
  +
  c A_{0,i}
  ,
  \label{eq:em-hamilton:dotAi}
  \\
  \Dot{p}^0 = -\fdv{H}{A_0} =
  - \pdv{H}{A_0}
  + \partial_{\alpha} \pdv{H}{A_{0,\alpha}}
  =
  -c p^{i}_{,i} + \frac{\sqrt{-g}}{c^2} j^0,
  \label{eq:em-hamilton:dotp0}
  \\
  \Dot{p}^i = - \fdv{H}{A_i} = - \pdv{H}{A_i}
  + \partial_{\alpha} \pdv{H}{A_{i,\alpha}}
  =
  - \frac{\sqrt{-g}}{c^2} j^{i}
  +
  \frac{\sqrt{-g}}{16\pi c} F^{\alpha i}_{,\alpha}
  .
  \label{eq:em-hamilton:dotpi}
\end{gather}

We show that the resulting system of equations is equivalent to Maxwell 
equations (which seems obvious).

From the equation~\eqref{eq:em-hamilton:dotAi} we obtain
\begin{equation}
  \label{eq:em-hamilton:dotAi:pi}
  p_i =
  \frac{\sqrt{-g}}{4\pi c^3}
  \qty(
  c A_{0,i} - \Dot{A}_{i} 
  )
  =
  \frac{\sqrt{-g}}{4\pi c^2}
  \qty(
  A_{0,i} - A_{i,0}
  )
  =
  \frac{\sqrt{-g}}{4\pi c^2}
  F_{i 0}.
\end{equation}
We raise the indices in the equation~\eqref{eq:em-hamilton:dotAi:pi} and substitute
the result in the equation~\eqref{eq:em-hamilton:dotpi}.
Writing the tensor components $F^{{\alpha}{\beta}}$
\begin{equation}
F^{{\alpha}{\beta}}=
\begin{pmatrix}
0 & -{D}^1 & -{D}^2 & -{D}^3 \\
{D}^1 & 0 & -{H}_3 & {H}_2 \\
{D}^2 & {H}_3 & 0 & -{H}_1 \\
{D}^3 & -{H}_2 & {H}_1 & 0
\end{pmatrix},
\label{eq:g^ab}
\end{equation}
we get the equation
\begin{equation}
  \frac{1}{\sqrt{{}^{3}g}}
  \qty[\partial_{{j}}H_{{k}}-\partial_{{k}}H_{{j}}]
  =
  -\frac{1}{c}\partial_t D^{{i}}+\frac{4\pi}{c}j^{{i}}.
\end{equation}

Similarly from~\eqref{eq:em-hamilton:dotp0} taking into account
equation~\eqref{eq:em-hamilton:dotAi:pi} and
connection~\eqref{eq:em-hamilton:varphi}
we obtain another Maxwell equation.
\begin{equation}
\frac{1}{\sqrt{{}^{3}g}}
\partial_{{i}} \qty(\sqrt{{}^{3}g} D^{{i}}) = 4\pi \rho.
\end{equation}

Thus, we showed that the resulting Hamiltonian gives
inhomogeneous Maxwell equations. It is obvious that the homogeneous Maxwell equations
is not obtained from the Hamiltonian and from the Bianchi identities for
tensor $F_{\alpha \beta}$.

\section{Conclusion}
\label{sec:conclusion}

Полученные уравнения
The authors applied the Dirac--Bergman method to the Maxwell equations. Obtained equations
is equivalent to inhomogeneous Maxwell equations. The main feature of our approach is that Maxwell's equations is considered in covariant form and in an arbitrary Riemannian coordinates.

\begin{acknowledgments}

The publication has been prepared with the support of the ``RUDN University Program 5-100''
and funded by Russian Foundation for Basic Research (RFBR) according to the research project
No~19-01-00645.

\end{acknowledgments}

 \bibliographystyle{elsarticle-num}

\bibliography{bib/hamilton-geom/cite}

\end{document}